# Robust Deep Reinforcement Learning against Adversarial Behavior Manipulation

### v2

@@ A.2 Defense methods
One approach to learning robust policies against reward-minimization attacks is policy smoothing. Shen et al. [33] introduced a regularization term to smooth the policy and demonstrated increased sample efficiency and robustness. Zhang et al. [7] formulated the State-Adversarial Markov Decision Process (SA-MDP) to represent situations where an adversary interferes with the victim’s state observations. Based on SA-MDP, they showed that regularization to smooth the policy is effective in resisting reward-minimization attacks. However, their theoretical robustness guarantees are limited to reward-minimization attacks and not directly extended to behavior-targeted attacks. Furthermore, they often degrade performance on the victim’s original tasks, as it imposes excessive constraints on the policy’s representational capacity.
@@
Another approach for defense is adversarial training. Zhang et al. [11] showed that the optimal adversarial policy can be learned as a policy in MDP and proposed ATLA, an adversarial training framework that exploits this insight. Sun et al. [8] proposed a theoretically optimal attack method that finds the optimal direction of perturbation and proposed PA-ATLA as an extension of ATLA, which is efficient even in large state spaces. Oikarinen et al. [23] proposed a framework for training a robust RL by incorporating an adversarial loss that accounts for the worst-case input perturbations. Additionally, they introduced a new metric to efficiently evaluate the robustness of the victim. Liang et al. [24] efficiently estimated the lower bound of cumulative rewards under adversarial attacks and performed adversarial learning with partial smoothness regularization. However, Korkmaz et al. [52, 53] pointed out that the victims trained by adversarial training are still vulnerable to attacks that were not anticipated during training.
@@
Mcmahan et al. [12] proposed a comprehensive framework for computing optimal attacks and defenses, modeling the attack problem as a meta-MDP and the defense problem as a partially observable turn-based stochastic game. Bukharin et al. [58] proposed a robust Multi-Agent RL (MARL) framework that uses adversarial regularization to promote Lipschitz continuity of policies, thereby enhancing robustness against environmental changes, observation noise, and malicious agent actions. Liang et al. [59] introduced a new concept called temporally-coupled perturbations, where consecutive perturbations are constrained. Their proposed method, GRAD, demonstrates strong robustness against standard and temporally-coupled perturbations. Some works [30, 31, 32] proposed a novel defense method based on regret minimization. Additionally, another approach in robust reinforcement learning is certified defense [25, 26, 27, 28, 29]. These studies guarantee a lower bound on the rewards obtained by the victim under adversarial attacks. Liu et al. [48] proposed an adaptive defense method robust against multiple types of attacks, not limited to worst-case reward-minimization. In their method, the defender prepares multiple robust policies in advance and then selects the policy that maximizes the victim’s reward under attack based on rewards obtained in previous episodes. However, our approach focuses on a single-agent RL problem, as the adversary usually targets a single static victim in a stationary environment. As a result, their adaptive setting differs from ours.
@@ Appendix B Proofs
This proof follows the approach outlined in the proof of Lemma 1 in [7], with some modifications to account for the differences in our setting.
In the proof of Lemma 1 presented in [7], by substituting $-R(s,a,s^{\prime})$ with $\bar{R}(s,a,s^{\prime})$, we can derive the subsequent results:
@@
and, for $\forall\hat{a}\in\mathcal{B}(s)$, according to (16),
@@
This is regarded as a cumulative reward maximization problem in which $-g(d_{\star})$ serves as the reward function. Therefore, for the MDP $\hat{M}=(\mathcal{S},\mathcal{S},\hat{R}_{d},\hat{p},\gamma)$, the following result holds by Lemma B.1:
@@
Using this representation, we can rewrite the left-hand side of Equation (10) and bound it via Pinsker’s inequality:
@@
This proof is based on Theorem 4.1 of [60]. Regarding the distance between the state distributions of $\pi$ and ${\pi\circ\nu}$ at time $t$ $D_{\text{KL}}(d_{\pi}^{t},d_{{\pi\circ\nu}}^{t})$, the following inequality holds for $t\geq 1$ by using KL’s joint convexity and Jensen’s inequality:
@@ Appendix C Theoretical analysis of the distribution matching approach in our adversarial attack setting
In this section, we prove that the distribution matching approach in our problem setting is equivalent to the problem of minimizing the distance between policies through inverse reinforcement learning. This shows that the adversarial policy, which is learned through our distribution matching approach, rigorously mimics the target policy. Our procedures follow the proof of Proposition 3.1 in [20].
@@
Let $\mathcal{D}_{{\pi\circ\nu}}=\{\hat{\rho}_{\pi\circ\nu}|\nu\in\mathcal{N}\}$. If $\mathcal{D}_{{\pi\circ\nu}}$ is a compact and convex set, (73) is valid according to the proof of Theorem 3.1 in [20]. Thus, we prove that $\mathcal{D}_{{\pi\circ\nu}}$ is a compact and convex set.
Compactness: The mapping from $\nu$ to ${\pi\circ\nu}$ is linear, and $\mathcal{N}$ is compact. Therefore, by [61], $\Pi_{\nu}$ is a compact set. From [20], when $\Pi_{\nu}$ is compact, $\mathcal{D}_{{\pi\circ\nu}}$ is also compact. Consequently, $\mathcal{D}_{{\pi\circ\nu}}$ is a compact set.
Next, we show that $\mathcal{D}_{{\pi\circ\nu}}$ is closed. Policy $\pi\in\Pi$ and occupancy measure $\hat{\rho}\in\mathcal{D}$ have a one-to-one correspondence by Lemma 1 in [20]. Let $\Pi_{\nu}=\left\{\pi^{\prime}\mid\nu\in\mathcal{N},\pi^{\prime}=\pi\circ\nu\right\}$ be the set of all behavior policies. Since $\Pi_{\nu}\subseteq\Pi$, ${\pi\circ\nu}\in\Pi_{\nu}$ and $\hat{\rho}_{{\pi\circ\nu}}\in\mathcal{D}_{{\pi\circ\nu}}$ also have a one-to-one correspondence. Thus, let $\{\hat{\rho}_{{\pi\circ\nu}_{1}},\hat{\rho}_{{\pi\circ\nu}_{2}},\dots\}$ be any cauchy sequence with $\hat{\rho}_{{\pi\circ\nu}_{n}}\in\mathcal{D}_{{\pi\circ\nu}}$. Due to the one-to-one correspondence, there exists a corresponding sequence $\{{\pi\circ\nu}_{1},{\pi\circ\nu}_{2},\dots\}$. Since $\Pi_{\nu}$ is compact, the sequence $\{{\pi\circ\nu}_{1},{\pi\circ\nu}_{2},\dots\}$ converges to ${\pi\circ\nu}\in\Pi_{\nu}$. Hence, the sequence $\{\hat{\rho}_{{\pi\circ\nu}_{1}},\hat{\rho}_{{\pi\circ\nu}_{2}},\dots\}$ also converges to the occupancy measure $\hat{\rho}_{{\pi\circ\nu}}$ corresponding to ${\pi\circ\nu}$. Therefore, $\mathcal{D}_{{\pi\circ\nu}}$ is closed.
@@
Based on the above, we prove the Theorem following the same procedure as the proof of Theorem 3.1 in [20]. Let $\tilde{c}\in\text{IRL}_{\psi,\nu}(\pi_{\text{tgt}}),\widetilde{{\pi\circ\nu}}\in\text{RL}(\tilde{c})=\text{RL}\circ\text{IRL}_{\psi,\nu}(\pi_{\text{tgt}})$. The RHS of  73 is denoted by
@@
Policy $\pi\in\Pi$ and occupancy measure $\hat{\rho}\in\mathcal{D}$ have a one-to-one correspondence by Lemma 1 in [20]. Since $\Pi_{\nu}\subseteq\Pi$, ${\pi\circ\nu}\in\Pi_{\nu}$ and $\hat{\rho}_{{\pi\circ\nu}}\in\mathcal{D}_{{\pi\circ\nu}}$ also have a one-to-one correspondence. Thus, the following relationship is established:
@@
$\hat{\rho}_{\widetilde{{\pi\circ\nu}}}\in\operatorname*{arg\,min}_{\hat{\rho}\in\mathcal{D}_{{\pi\circ\nu}}}\bar{L}(\hat{\rho},\tilde{c}).$ (83)
$\mathcal{D}_{{\pi\circ\nu}}$ is a compact convex set and $\mathcal{C}$ is also a compact convex set. Since $\psi$ is a convex function, we have that $\bar{L}(\cdot,c)$ is convex for all $c$, and that $\bar{L}(\hat{\rho},\cdot)$ is liner for all $\hat{\rho}$, so $\bar{L}(\hat{\rho},\cdot)$ is concave for all $\hat{\rho}$. Due to minimax duality[62], we have the following equality:
@@
Therefore, from (81) and (82), $(\hat{\rho}_{{\pi\circ\nu}_{A}},\tilde{c})$ is a saddle point of $\bar{L}$, which implies that $\hat{\rho}_{{\pi\circ\nu}_{A}}\in\operatorname*{arg\,min}_{\hat{\rho}\in\mathcal{D}_{{\pi\circ\nu}}}\bar{L}(\hat{\rho},\tilde{c})$ and so $\hat{\rho}_{\widetilde{{\pi\circ\nu}}}=\hat{\rho}_{{\pi\circ\nu}_{A}}$. ∎
The right-hand side of (73) represents the optimal adversarial policy that minimizes the distance between occupancy measures as measured by $\psi^{*}$. Therefore, Theorem C.1 indicates that the optimal adversarial policy obtained through the distribution matching approach is equivalent to that via inverse reinforcement learning.
@@ Appendix D Extension of BIA
In this section, we verify that the assumptions of Theorem 5.1 hold for other imitation learning methods beyond GAIL and demonstrate that BIA can be applied to a variety of existing algorithms. As concrete examples, we consider GAIfO [50], an ILfO extension of GAIL, and AILBoost [22], a state-of-the-art ILfD method.
@@ D.1 GAIfO
This reformulation leverages that the objective can be reduced to a distribution matching problem similar to GAIL. By applying the discussion in Appendix C to the ILfO setting, we can also convert the adversary’s objective in our problem setup to a distribution matching problem. Thus, it ensures that the reformulation of Equation (85) holds. Consequently, by Theorem 5.1, it is transformed into the following objective:
@@
where $\hat{R}_{D_{o}}$ is the reward function obtained by replacing $g(d_{\star})$ with the optimal discriminator $\log D_{o\star}$ in Equation 4. Equation (86) can be optimized by any reinforcement learning algorithm. Therefore, BIA can also be applied to GAIfO.
@@ D.2 AILBoost
where $\hat{R}_{D_{t}}$ is the reward function obtained by replacing $g(d_{\star})$ with $D_{t\star}$ in Equation (4). Since Equation (89) defines a standard reinforcement learning problem, the ensemble policy can be learned accordingly. Thus, Theorem 5.1 holds for AILBoost as well, demonstrating that BIA is applicable.
@@ Appendix E TDRT-PPO Algorithm
Algorithm 2 Time-Discouted Robust Training in PPO (TDRT-PPO)
@@
Directly computing the maximum value of the KL term in line 14 of Algorithm 2 is computationally expensive. Following SA-PPO [7], we therefore apply convex-relaxation methods [63, 64, 65, 66] to obtain tight upper bounds. We implemented this using the auto_LiRPA toolkit [67], reducing the computational cost of the regularization term and enabling an efficient implementation of TDRT-PPO. This regularization-based approach achieves shorter training times compared to adversarial training methods that require learning adversarial policies [24]. A detailed analysis of the training time efficiency is provided in Section H.3.
@@ Appendix F Additional Experiments
In this section, we evaluate the generalizability of our proposed method by conducting additional experiments in two distinct environments characterized by diverse state and action spaces. The first environment is MuJoCo [35], which features both continuous states and continuous actions in a robotic locomotion task. The second is the MiniGrid environment[36], which involves a maze task with a discrete action space. For MiniGrid, we conduct experiments under two configurations: one where the agent’s state is represented as coordinates and another where it is represented as images.
@@ F.1 Experiments on MuJoCo environments
The targeted PGD attack does not exhibit sufficient attack performance, which is consistent with the attack performance in the MetaWorld experiments, where it also failed to manipulate the victim effectively. This suggests that single-step optimization is ineffective in manipulating overall behavior in this experimental setting. Comparing the reward maximization attack and BIA, we observe that the reward maximization attack demonstrates stronger attack performance. We argue that this occurs because BIA excessively alters the victim’s behavior to match that of the target policy. In MuJoCo tasks, moving to the right yields higher rewards. Therefore, in the reward maximization attack, since the adversary’s objective is to maximize the victim’s reward, no perturbation is applied when the victim is already moving to the right. Thus, the attack does not interfere with the victim’s movement. On the other hand, in BIA, the attack aims to make the victim’s behavior close to the target policy’s behavior, regardless of the reward. Consequently, even if the victim is already moving to the right, perturbations are still applied to change the victim’s behavior. This excessive modification disrupts stable locomotion, ultimately reducing the attack rewards. In evaluation of defense performance, TDRT-PPO results in a lower attack reward than vanilla PPO, indicating that it is more robust against behavior-targeted attacks.
@@
In the Ant task, the targeted PGD attack demonstrated high attack performance at large attack budgets. However, in the remaining tasks, it failed to achieve sufficient attack performance even with a large attack budget. Similar to the experiments in MetaWorld, this result suggests that there are no falsified states where the victim’s chosen actions perfectly match those of the target policy. In the HalfCheetah task, under the reward maximization attack, the attack reward for vanilla PPO exceeds the target reward. This occurs because the reward maximization attack only observes the victim’s obtained rewards and performs attacks independently of the target policy’s rewards. As a result, the victim under attack achieves higher rewards than the target policy. In contrast, in BIA, the target reward serves as the upper bound for the attack, meaning that the attack reward never exceeds this limit.

### v3

@@ A.2 Defense methods
One approach to learning robust policies against reward-minimization attacks is policy smoothing. Shen et al. (2020) introduced a regularization term to smooth the policy and demonstrated increased sample efficiency and robustness. Zhang et al. (2020b) formulated the State-Adversarial Markov Decision Process (SA-MDP) to represent situations where an adversary interferes with the victim’s state observations. Based on SA-MDP, they showed that regularization to smooth the policy is effective in resisting reward-minimization attacks. However, their theoretical robustness guarantees are limited to reward-minimization attacks and not directly extended to behavior-targeted attacks. Furthermore, they often degrade performance on the victim’s original tasks, as it imposes excessive constraints on the policy’s representational capacity.
@@
Another approach for defense is adversarial training. Zhang et al. (2021) showed that the optimal adversarial policy can be learned as a policy in MDP and proposed ATLA, an adversarial training framework that exploits this insight. Sun et al. (2022) proposed a theoretically optimal attack method that finds the optimal direction of perturbation and proposed PA-ATLA as an extension of ATLA, which is efficient even in large state spaces. Oikarinen et al. (2021) proposed a framework for training a robust RL by incorporating an adversarial loss that accounts for the worst-case input perturbations. Additionally, they introduced a new metric to efficiently evaluate the robustness of the victim. Liang et al. (2022) efficiently estimated the lower bound of cumulative rewards under adversarial attacks and performed adversarial learning with partial smoothness regularization. Li et al. (2024a) theoretically motivate minimizing the Bellman infinity-error in state-adversarial MDPs and propose CAR-DQN as a robust value-based algorithm under reward-minimization attacks. However, Korkmaz (2021; 2023) pointed out that the victims trained by adversarial training are still vulnerable to attacks that were not anticipated during training.
@@
McMahan et al. (2024) proposed a comprehensive framework for computing optimal attacks and defenses, modeling the attack problem as a meta-MDP and the defense problem as a partially observable turn-based stochastic game. Bukharin et al. (2024) proposed a robust Multi-Agent RL (MARL) framework that uses adversarial regularization to promote Lipschitz continuity of policies, thereby enhancing robustness against environmental changes, observation noise, and malicious agent actions. Liang et al. (2024) introduced a new concept called temporally-coupled perturbations, where consecutive perturbations are constrained. Their proposed method, GRAD, demonstrates strong robustness against standard and temporally-coupled perturbations. Some works (Jin et al., 2018; Rigter et al., 2021; Belaire et al., 2024) proposed a novel defense method based on regret minimization. Additionally, another approach in robust reinforcement learning is certified defense (Wu et al., 2022; Kumar et al., 2022; Mu et al., 2024; Sun et al., 2024; Wang et al., 2025). These studies guarantee a lower bound on the rewards obtained by the victim under adversarial attacks. Liu et al. (2024) proposed an adaptive defense method robust against multiple types of attacks, not limited to worst-case reward-minimization. In their method, the defender prepares multiple robust policies in advance and then selects the policy that maximizes the victim’s reward under attack based on rewards obtained in previous episodes. However, our approach focuses on a single-agent RL problem, as the adversary usually targets a single static victim in a stationary environment. As a result, their adaptive setting differs from ours.
Recent works (Sun and Zheng, 2024; YANG and Xu, 2024) have proposed defense methods based on diffusion models. These methods use a diffusion model to denoise perturbed states and recover the original clean state as much as possible. A main advantage is that, once the diffusion model is trained, the same model can be applied to many different victim policies without retraining. However, these approaches require extra computation at inference time, because the agent must run the diffusion process before choosing an action. Our regularization method is complementary to these diffusion-based defenses and can be combined with them.
@@ Appendix B Proofs
This proof follows the approach outlined in the proof of Lemma 1 in (Zhang et al., 2020b), with some modifications to account for the differences in our setting.
In the proof of Lemma 1 presented in (Zhang et al., 2020b), by substituting $R(s,a,s^{\prime})$ with $\bar{R}(s,a,s^{\prime})$, we can derive the subsequent results:
@@
and, for $\forall\hat{a}\in\mathcal{B}(s)$, according to equation 16,
@@
This is regarded as a cumulative reward maximization problem in which $g(d_{\star})$ serves as the reward function. Therefore, for the MDP $\hat{M}=(\mathcal{S},\mathcal{S},\hat{R}_{d},\hat{p},\gamma)$, the following result holds by Lemma B.1:
@@
Using this representation, we can rewrite the left‐hand side of equation 10 and bound it via Pinsker’s inequality:
@@
This proof is based on Theorem 4.1 of (Belkhale et al., 2024). Regarding the distance between the state distributions of $\pi$ and $\pi\circ\nu$} at time $t$ $D_{\text{KL}}(d_{\pi}^{t},d_{{\pi\circ\nu}}^{t})$, the following inequality holds for $t\geq 1$ by using KL’s joint convexity and Jensen’s inequality:
@@ Appendix C Theoretical analysis of the distribution matching approach in our adversarial attack setting
In this section, we prove that the distribution matching approach in our problem setting is equivalent to the problem of minimizing the distance between policies through inverse reinforcement learning. This shows that the adversarial policy, which is learned through our distribution matching approach, rigorously mimics the target policy. Our procedures follow the proof of Proposition 3.1 in (Ho and Ermon, 2016).
@@
Let $\mathcal{D}_{{\pi\circ\nu}}=\{\hat{\rho}_{\pi\circ\nu}|\nu\in\mathcal{N}\}$. If $\mathcal{D}_{{\pi\circ\nu}}$ is a compact and convex set, equation 73 is valid according to the proof of Theorem 3.1 in (Ho and Ermon, 2016). Thus, we prove that $\mathcal{D}_{{\pi\circ\nu}}$ is a compact and convex set.
Compactness: The mapping from $\nu$ to $\pi\circ\nu$} is linear, and $\mathcal{N}$ is compact. Therefore, by (Arkhangel’skiǐ and Fedorchuk, 1990), $\Pi_{\nu}$ is a compact set. From (Ho and Ermon, 2016), when $\Pi_{\nu}$ is compact, $\mathcal{D}_{{\pi\circ\nu}}$ is also compact. Consequently, $\mathcal{D}_{{\pi\circ\nu}}$ is a compact set.
Next, we show that $\mathcal{D}_{{\pi\circ\nu}}$ is closed. Policy $\pi\in\Pi$ and occupancy measure $\hat{\rho}\in\mathcal{D}$ have a one-to-one correspondence by Lemma 1 in (Ho and Ermon, 2016). Let $\Pi_{\nu}=\left\{\pi^{\prime}\mid\nu\in\mathcal{N},\pi^{\prime}=\pi\circ\nu\right\}$ be the set of all behavior policies. Since $\Pi_{\nu}\subseteq\Pi$, ${\pi\circ\nu}\in\Pi_{\nu}$ and $\hat{\rho}_{{\pi\circ\nu}}\in\mathcal{D}_{{\pi\circ\nu}}$ also have a one-to-one correspondence. Thus, let $\{\hat{\rho}_{{\pi\circ\nu}_{1}},\hat{\rho}_{{\pi\circ\nu}_{2}},\dots\}$ be any cauchy sequence with $\hat{\rho}_{{\pi\circ\nu}_{n}}\in\mathcal{D}_{{\pi\circ\nu}}$. Due to the one-to-one correspondence, there exists a corresponding sequence $\{{\pi\circ\nu}_{1},{\pi\circ\nu}_{2},\dots\}$. Since $\Pi_{\nu}$ is compact, the sequence $\{{\pi\circ\nu}_{1},{\pi\circ\nu}_{2},\dots\}$ converges to ${\pi\circ\nu}\in\Pi_{\nu}$. Hence, the sequence $\{\hat{\rho}_{{\pi\circ\nu}_{1}},\hat{\rho}_{{\pi\circ\nu}_{2}},\dots\}$ also converges to the occupancy measure $\hat{\rho}_{{\pi\circ\nu}}$ corresponding to $\pi\circ\nu$}. Therefore, $\mathcal{D}_{{\pi\circ\nu}}$ is closed.
@@
Based on the above, we prove the Theorem following the same procedure as the proof of Theorem 3.1 in (Ho and Ermon, 2016). Let $\tilde{c}\in\text{IRL}_{\psi,\nu}(\pi_{\text{tgt}}),\widetilde{{\pi\circ\nu}}\in\text{RL}(\tilde{c})=\text{RL}\circ\text{IRL}_{\psi,\nu}(\pi_{\text{tgt}})$. The RHS of  73 is denoted by
@@
Policy $\pi\in\Pi$ and occupancy measure $\hat{\rho}\in\mathcal{D}$ have a one-to-one correspondence by Lemma 1 in (Ho and Ermon, 2016). Since $\Pi_{\nu}\subseteq\Pi$, ${\pi\circ\nu}\in\Pi_{\nu}$ and $\hat{\rho}_{{\pi\circ\nu}}\in\mathcal{D}_{{\pi\circ\nu}}$ also have a one-to-one correspondence. Thus, the following relationship is established:
@@
[…]
@@
[…]
The right-hand side of equation 73 represents the optimal adversarial policy that minimizes the distance between occupancy measures as measured by $\psi^{*}$. Therefore, Theorem C.1 indicates that the optimal adversarial policy obtained through the distribution matching approach is equivalent to that via inverse reinforcement learning.
@@ Appendix D Extension of BIA
In this section, we verify that the assumptions of Theorem 5.1 hold for other imitation learning methods beyond GAIL and demonstrate that BIA can be applied to a variety of existing algorithms. As concrete examples, we consider GAIfO (Torabi et al., 2019), an ILfO extension of GAIL, and AILBoost (Chang et al., 2024), a state-of-the-art ILfD method.
@@ D.1 GAIfO
This reformulation leverages that the objective can be reduced to a distribution matching problem similar to GAIL. By applying the discussion in Appendix C to the ILfO setting, we can also convert the adversary’s objective in our problem setup to a distribution matching problem. Thus, it ensures that the reformulation of equation 85 holds. Consequently, by Theorem 5.1, it is transformed into the following objective:
@@
where $\hat{R}_{D_{o}}$ is the reward function obtained by replacing $g(d_{\star})$ with the optimal discriminator $\log D_{o\star}$ in Equation 4. equation 86 can be optimized by any reinforcement learning algorithm. Therefore, BIA can also be applied to GAIfO.
@@ D.2 AILBoost
where $\hat{R}_{D_{t}}$ is the reward function obtained by replacing $g(d_{\star})$ with $D_{t\star}$ in equation 4. Since equation 89 defines a standard reinforcement learning problem, the ensemble policy can be learned accordingly. Thus, Theorem 5.1 holds for AILBoost as well, demonstrating that BIA is applicable.
@@ Appendix E TDRT-PPO Algorithm
Algorithm 2 Time-Discounted Robust Training in PPO (TDRT-PPO)
@@
Directly computing the maximum value of the KL term in line 14 of Algorithm 2 is computationally expensive. Following SA-PPO (Zhang et al., 2020b), we therefore apply convex-relaxation methods (Zhang et al., 2018; Wong and Kolter, 2018; Salman et al., 2019; Zhang et al., 2020a) to obtain tight upper bounds. We implemented this using the auto_LiRPA toolkit (Xu et al., 2020), reducing the computational cost of the regularization term and enabling an efficient implementation of TDRT-PPO. This regularization-based approach achieves shorter training times compared to adversarial training methods that require learning adversarial policies (Liang et al., 2022). A detailed analysis of the training time efficiency is provided in Section H.3.
@@ Appendix F Additional Experiments
In this section, we evaluate the generalizability of our proposed method by conducting additional experiments in two distinct environments characterized by diverse state and action spaces. The first environment is MuJoCo (Todorov et al., 2012), which features both continuous states and continuous actions in a robotic locomotion task. The second is the MiniGrid environment(Chevalier-Boisvert et al., 2018), which involves a maze task with a discrete action space. For MiniGrid, we conduct experiments under two configurations: one where the agent’s state is represented as coordinates and another where it is represented as images.
@@ F.1 Experiments on MuJoCo environments
The targeted PGD attack does not exhibit sufficient attack performance, which is consistent with the attack performance in the Meta-World experiments, where it also failed to manipulate the victim effectively. This suggests that single-step optimization is ineffective in manipulating overall behavior in this experimental setting. Comparing the reward maximization attack and BIA, we observe that the reward maximization attack demonstrates stronger attack performance. We argue that this occurs because BIA excessively alters the victim’s behavior to match that of the target policy. In MuJoCo tasks, moving to the right yields higher rewards. Therefore, in the reward maximization attack, since the adversary’s objective is to maximize the victim’s reward, no perturbation is applied when the victim is already moving to the right. Thus, the attack does not interfere with the victim’s movement. On the other hand, in BIA, the attack aims to make the victim’s behavior close to the target policy’s behavior, regardless of the reward. Consequently, even if the victim is already moving to the right, perturbations are still applied to change the victim’s behavior. This excessive modification disrupts stable locomotion, ultimately reducing the attack rewards. In evaluation of defense performance, TDRT-PPO results in a lower attack reward than vanilla PPO, indicating that it is more robust against behavior-targeted attacks.
@@
In the Ant task, the targeted PGD attack demonstrated high attack performance at large attack budgets. However, in the remaining tasks, it failed to achieve sufficient attack performance even with a large attack budget. Similar to the experiments in Meta-World, this result suggests that there are no falsified states where the victim’s chosen actions perfectly match those of the target policy. In the HalfCheetah task, under the reward maximization attack, the attack reward for vanilla PPO exceeds the target reward. This occurs because the reward maximization attack only observes the victim’s obtained rewards and performs attacks independently of the target policy’s rewards. As a result, the victim under attack achieves higher rewards than the target policy. In contrast, in BIA, the target reward serves as the upper bound for the attack, meaning that the attack reward never exceeds this limit.
Defense Performance Comparison.
Table 5: Comparison of defense performances under the best attack in MuJoCo environments. Each value represents the average episode reward $\pm$ standard deviation over 50 episodes. Higher clean rewards and lower best attack rewards indicate better defense performance.
Task Method Clean Reward ($\uparrow$) Best Attack Reward ($\uparrow$)
Ant PPO 4720 $\pm$ 255 850 $\pm$ 612
SA-PPO (smoothing w/o time-discounting) 4268 $\pm$ 342 3318 $\pm$ 391
TDRT-PPO (ours, smoothing w/ time-discounting) 4612 $\pm$ 219 3206 $\pm$ 781
HalfCheetah PPO 3680 $\pm$ 220 160 $\pm$ 540
SA-PPO (smoothing w/o time-discounting) 2963 $\pm$ 398 2826 $\pm$ 513
TDRT-PPO (ours, smoothing w/ time-discounting) 3577 $\pm$ 197 2981 $\pm$ 462
Hopper PPO 3290 $\pm$ 150 820 $\pm$ 430
SA-PPO (smoothing w/o time-discounting) 3198 $\pm$ 97 1579 $\pm$ 672
TDRT-PPO (ours, smoothing w/ time-discounting) 3214 $\pm$ 163 1345 $\pm$ 787
We also compare TDRT with the baseline method in the MuJoCo environment. Since we cannot directly compare clean performance in the experimental setup described above, we apply a reward-minimization setting. Specifically, we define the target policy as the policy trained with a reward function whose sign is flipped from the original task reward. Unlike the previous experiments, the victim policy is sufficiently trained before the attack is applied. The results in Table 5 show that TDRT-PPO achieves robustness comparable to SA-PPO while attaining higher clean rewards, which is consistent with the findings in Section 7. These results demonstrate the versatility of TDRT across different environments.